A manufacturer of ball pens claims that a certain pen they manufactures has a mean writing life of 400 pages. A purchasing agent selects a sample of 100 pens and puts them for test. The mean writing life from the sample was 390 pages with a standard deviation of 30 at alpha 0.01.


1
Expert's answer
2022-05-31T11:21:32-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=400"

"H_1:\mu\not=400"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=99" and the critical value for a two-tailed test is "t_c = 2.626405."

The rejection region for this two-tailed test is "R = \{t:|t|> 2.626405\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{390-400}{30/\sqrt{100}}=-3.3333"

Since it is observed that "|t|=3.3333> 2.626405=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=99" degrees of freedom, "t=-3.3333" is "p=0.001208," and since "p=0.001208<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 400, at the "\alpha = 0.01" significance level.